OpenStudy (anonymous):

I need to find an equation having the given slope and containing the given point and then put it in y=mx+b form. m= -6, (4,0)

OpenStudy (amistre64):

0 = -6(4) +b

OpenStudy (anonymous):

y = -6x +b Now plug in the coordinates 0 = -24 +b; b = 24 y = -6x +24

OpenStudy (anonymous):

You are great at explaining this to me.. Thanks